# Mohamed Bouadi, Pratinav Seth Aditya Tanna, Vinay Kumar Sankarapu Lexsi Labs, India & France

### v1

@@ 1  Introduction
Tabular data remain the most prevalent form of data in real-world applications, spanning critical systems across healthcare, finance, and scientific research. Despite the remarkable progress of deep learning in natural language processing [27, 42] and computer vision [11], gradient boosted trees (GBTs) remain the predominant state-of-the-art (SOTA) for tabular prediction tasks. In other data modalities, foundation models—particularly Large Language Models (LLMs) [46, 26]—have significantly advanced the ability to tackle new tasks and few-shot learning. This is largely due to their remarkable in-context learning (ICL) capabilities [45, 4], which enable them to capture patterns directly from prompts without updating their parameters. This success combined with the pervasiveness of tables have spurred interest in tabular foundation models [38].
Although LLMs are primarily designed to process natural language, recent efforts have explored fine-tuning them for tabular data tasks [14, 8]. These approaches typically rely on table serialization, which is the process of converting table rows into text or sentences suitable for tokenization. For instance, [9] fine-tuned a Llama 3-8B model on a large corpus of serialized tables and demonstrated that this strategy can outperform traditional tree-based models in few-shot scenarios. However, such language model–based approaches face inherent challenges. Their limited context windows restrict the number of serialized examples that can be processed simultaneously (e.g., up to 32 or 64 shots in [9]), and it remains uncertain whether LLMs can reliably interpret and reason over numerical values [37].
Adopting a fundamentally different strategy, the authors of [16] introduced TabPFN, a transformer-based tabular foundation model designed for classification tasks and pretrained exclusively on synthetic tabular data. A key feature of TabPFN is its ability to perform in-context learning directly on tables, removing the need for tokenization and allowing efficient processing of relatively small datasets—up to 1K samples and 100 features. Building on this foundation, TabICL [32] introduced a simplified three-component architecture comprising: (1) column-wise embeddings via Set Transformers to capture distribution-aware feature semantics, (2) row-wise interactions with rotary positional encodings to model inter-feature dependencies, and (3) dataset-level ICL prediction through split attention, ensuring a clear separation between training and test samples. These developments position tabular foundation models as a compelling alternative to traditional approaches, particularly for zero-shot prediction tasks where dataset-specific training is infeasible.
@@
The column-wise embedding component of Orion-MSP follows TabICL’s approach [32], using Set Transformers with Induced Set Attention Blocks (ISAB) [25] to create distribution-aware feature embeddings in a permutation-invariant manner. The multi-scale row interaction component processes these embeddings at multiple resolutions, with each scale using sparse attention patterns tailored to its granularity. The resulting multi-scale representations are aggregated into unified row embeddings, which then interact with the Perceiver memory before proceeding to the final ICL prediction stage. This ICL component employs split attention with label injection, ensuring proper train-test separation.
@@ 2  Related Work
Tabular In-Context Learning: The application of in-context learning (ICL) to tabular data has recently attracted significant attention. TabPFN [16] pioneered this direction by meta-training a transformer on synthetically generated datasets using structural causal models. Its encoder–decoder design allows test samples to attend to training examples, enabling zero-shot predictions without gradient-based fine-tuning. While TabPFN demonstrated strong performance on small datasets, its alternating column- and row-wise attention mechanisms make scaling to larger tables computationally prohibitive.
TabDPT [28] showed that comparable performance can be achieved on real-world datasets by using similarity-based retrieval to construct contextual examples—an idea first explored in TabR [12]. The authors extended this paradigm by integrating diffusion-based representation learning, improving robustness to missing values and distributional shifts. However, the diffusion process introduces substantial computational overhead and retains dense attention, limiting scalability. Similarly, TabPFN-v2 [17] introduced cell-based in-context learning, extending row-wise encoding to datasets exceeding 10,00010{,}000 samples, but it still inherits quadratic attention costs in high-dimensional tables.
Building on these foundations, TabICL [32] proposed a table-native transformer architecture with three components: column embedding via Set Transformers, row-wise interaction with rotary positional encodings, and an in-context learning prediction module. This design achieved state-of-the-art results across diverse benchmarks while maintaining architectural simplicity and training efficiency. Nonetheless, dense attention in row interactions and the strictly sequential pipeline limit iterative refinement, cross-component communication, and scalability to tables with more than 100 features.
ContextTab [35] further enhanced tabular in-context learning by incorporating contextualized feature embeddings and attention mechanisms tailored for heterogeneous tabular data. While improving performance in complex datasets, it still processes features at a single scale and relies on dense attention, limiting computational efficiency on high-dimensional tables.
@@
Sparse Attention Mechanisms: Sparse attention techniques from natural language processing offer a promising route to improve computational efficiency in tabular in-context learning. BigBird [43] and Longformer [1] demonstrated that block-sparse attention patterns can approximate dense attention with linear complexity while maintaining strong theoretical guarantees. Similarly, Sparse Transformers [5, 20] employ structured sparsity for generative modeling, reducing computation without substantial performance degradation. Despite their success in sequential data, these methods have yet to be systematically adapted for tabular in-context learning, where the primary challenge lies in feature dimension rather than sequence length.
Hierarchical and Multi-Scale Architectures: Hierarchical architectures have proven effective in other domains. Funnel Transformers [6] and Swin Transformers [30] use multi-scale processing and pooling to capture information at different resolutions, while Set Transformers [34, 10] leverage pooling by multihead attention for permutation-invariant set processing. Although TabICL [32] employs Set Transformers for column embeddings, it does not incorporate hierarchical multi-scale processing or iterative pooling across feature groups, limiting its ability to model complex interactions in high-dimensional tables.
Cross-Component Communication: Cross-component memory and iterative refinement have shown success in multimodal learning. Perceiver [19] and Perceiver IO [18] introduce latent bottlenecks to compress and share information across modalities, and vision-language models [39] leverage iterative cross-attention for refinement. However, these approaches do not address the causal constraints of in-context learning, where test examples must never influence the representation of training data, leaving a gap for tabular in-context learning.
@@ 3.2  High-level Structure: From Data to ICL
Orion-MSP extends the original TabICL [32] architecture with three complementary innovations designed to address the fundamental challenges of tabular data processing: computational inefficiency, limited feature interaction modeling, and the need for hierarchical pattern recognition. Our approach maintains the core in-context learning paradigm while introducing architectural enhancements that significantly improve both efficiency and performance.
@@ 3.3  Column-wise Embedding
Tabular data exhibits unique characteristics compared to other modalities: each column represents a distinct feature with its own distribution, scale, and statistical properties (e.g., mean, variance, skewness, kurtosis). To capture these distributional characteristics, we adopt the original TabICL [32] column-wise embedder to map each scalar cell in a column $c_{j}\in\mathbb{R}^{n}$ to a $d$-dimensional representation using a shareable Set Transformer, $TF_{\text{col}}$, that treats the column as a permutation-invariant set of values. Our goal is to transform each cell value $X_{ij}$ into a $d$-dimensional embedding $\mathbf{E}_{ij}\in\mathbb{R}^{d}$ that encodes both:
@@ 3.3.2  Induced Set Attention Blocks (ISAB):
To efficiently capture global distributional information while maintaining computational tractability, we employ ISAB [25] with $k$ learnable inducing points. It consists of two sequential Multi-Head Attention Blocks $(\text{MAB}_{1},\text{MAB}_{2})$:
@@
Following TabICL [32], we use d=128d{=}128, k=128k{=}128, 4 heads, and 3 ISAB blocks. Crucially, in Equation 2, we use only training samples $\mathbf{U}_{j}^{\text{train}}\in\mathbb{R}^{n_{\text{train}}\times d}$ as keys and values. This ensures that the inducing points $\mathbf{M}_{j}$ capture the distribution of the training data only, preventing information leakage from test samples during embedding. This is crucial for maintaining the in-context learning paradigm.
@@ 3.4.1  Motivation and Design Principles
Inspired by hierarchical representations in vision [41] and multi-resolution modeling in speech [44], Orion-MSP decomposes feature interactions into multiple resolution levels:
@@
To further improve efficiency and generalization, we adopt a block-sparse attention pattern inspired by Longformer [1] and BigBird [43], as depicted in Figure 2:
@@ 3.4.2  Multi-Scale Feature Grouping:
The default grouping strategy uses a learnable soft grouping via Pooling by Multihead Attention (PMA) [24] to adaptively attend to features:
@@ 3.4.5  Transformer Encoder per Scale:
  • Rotary positional encoding (RoPE): Applied to queries and keys before attention [36]
@@ 3.5  Cross-Component Memory with Perceiver Architecture
To overcome this limitation, we introduce a Perceiver-style latent memory module [19] that enables safe, leak-free communication between architectural components. This latent memory acts as a shared representation space that can be written to by training samples and read from by both training and test samples, ensuring compliance with ICL constraints while promoting global knowledge sharing.
In standard transformer-based tabular architectures such as TabICL [32], model components operate in a strictly sequential and isolated fashion:
@@
Inspired by the Perceiver architecture [19], we introduce a learnable latent memory $\mathbf{L}\in\mathbb{R}^{P\times d}$ with $P$ memory slots. The key idea is:
@@ 4.1  Experimental Setting
Our experimental evaluation spans three widely recognized benchmark suites: TALENT [40] (181 automatically discovered classification datasets), OpenML-CC18 [2] (72 curated datasets), and TabZilla [29] (36 heterogeneous tasks). Together, these benchmarks enable a comprehensive assessment across diverse tabular learning scenarios. In addition, we perform domain-specific evaluations in high-impact application areas such as healthcare and finance to examine the real-world relevance of our method. All experiments strictly follow the official dataset splits provided by each benchmark to ensure reproducibility and fairness.
@@
Finally, we emphasize that models with higher mean ranks may not always achieve the highest absolute accuracy or F1-scores on every dataset. Rankings are computed per dataset and then averaged across all datasets, providing a normalized indicator of overall consistency rather than peak task-specific performance. In contrast, absolute metrics highlight maximum achievable performance on individual tasks. Comprehensive dataset statistics are presented in Appendix A.4.
@@
We compare our model with six state-of-the-art tabular foundation models: TabPFN [16], TabICL [32], OrionBiX, Mitra, ContextTab [35], and TabDPT [28]. In addition, we include established traditional baselines using autogloun [7] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.
@@
Performance. We measure predictive capability using standard classification metrics—Accuracy (ACC), AUC-ROC, and weighted F1-score (F1)—computed across the benchmark suites TALENT, OpenML-CC18, and TabZilla. These benchmarks encompass datasets with diverse characteristics, including varying sample sizes, feature dimensionalities, and class balance, allowing a comprehensive assessment of model generalization.
@@ 4.2  Results
Table 1: Performance comparison across hree benchmark suites—TALENT, OpenML-CC18, and TabZilla. Ranks denote the mean rank per benchmark suite (lower is better). Metrics: ACC = Accuracy, F1 = Weighted F1. The “All” column reports the aggregated rank across all suites. Formatting: 1st place; 2nd place.
@@
Table 2: Performance variation by dataset size across all benchmark suites. ACC = Accuracy; F1 = Weighted F1-score, averaged across datasets within each size category: Small (<<1K samples), Medium (1K–10K), and Large (>>10K).
@@
ACC F1 ACC F1 ACC F1
XGBoost 0.8168 0.7964 0.8363 0.8314 0.8969 0.8920
CatBoost 0.8124 0.7935 0.8340 0.8264 0.8797 0.8733
Random Forest 0.7988 0.8187 0.8285 0.8221 0.8694 0.8628
LightGBM 0.8143 0.7789 0.8314 0.8226 0.8827 0.8764
TabICL 0.8301 0.8338 0.8486 0.8398 0.8802 0.8743
OrionBiX 0.8330 0.8150 0.8348 0.8260 0.8729 0.8670
OrionMSP 0.8232 0.8194 0.8494 0.8402 0.8843 0.8768
TabPFN 0.8325 0.8131 0.8557 0.8462 0.8783 0.8713
Mitra 0.4334 0.3236 0.3600 0.2553 0.3837 0.2754
ContextTab 0.7578 0.7363 0.6210 0.5566 0.6388 0.5638
TabDPT 0.8333 0.8271 0.8424 0.8339 0.8831 0.8765
@@
Table 3: Performance variation by feature dimensionality (dataset width) across all benchmark suites. ACC = Accuracy; F1 = Weighted F1-score, averaged across datasets within each width category. Values are on a 0–1 scale (higher is better). Models are grouped by adaptation strategy. Formatting: 1st place ; 2nd place within each group.
@@
ACC F1 ACC F1 ACC F1
XGBoost 0.8222 0.8159 0.8482 0.8410 0.9140 0.9039
CatBoost 0.8145 0.8067 0.8441 0.8344 0.9157 0.9084
Random Forest 0.8005 0.7044 0.8410 0.8235 0.9034 0.8936
LightGBM 0.8128 0.7907 0.8458 0.8326 0.8999 0.8908
TabICL 0.8208 0.8119 0.8627 0.8549 0.9101 0.8936
OrionBiX 0.8112 0.8043 0.8510 0.8417 0.8859 0.8849
OrionMSP 0.8394 0.8314 0.8572 0.8478 0.8860 0.8837
TabPFN 0.8187 0.8092 0.8676 0.8589 0.9129 0.9111
Mitra 0.3737 0.2683 0.3886 0.2781 0.2521 0.1497
ContextTab 0.6391 0.5719 0.6480 0.5843 0.6017 0.5651
TabDPT 0.8262 0.8189 0.8566 0.8483 0.8845 0.8820

### v3

@@ 1  Introduction
Tabular data remain the most prevalent form of data in real-world applications, spanning critical systems across healthcare, finance, and scientific research. Despite the remarkable progress of deep learning in natural language processing [liu2023pre, yu2024natural] and computer vision [goldblum2023battle], gradient boosted trees (GBTs) remain the predominant state-of-the-art (SOTA) for tabular prediction tasks. In other data modalities, foundation models—particularly Large Language Models (LLMs) [DBLP:conf/nips/00020D24, DBLP:conf/nips/LiZLML0HY24]—have significantly advanced the ability to tackle new tasks and few-shot learning. This is largely due to their remarkable in-context learning (ICL) capabilities [DBLP:conf/iclr/Zhang0YOKZY025, DBLP:conf/uss/CarliniTWJHLRBS21], which enable them to capture patterns directly from prompts without updating their parameters. This success combined with the pervasiveness of tables have spurred interest in tabular foundation models [DBLP:conf/icml/BreugelS24].
Although LLMs are primarily designed to process natural language, recent efforts have explored fine-tuning them for tabular data tasks [tabllm, tmlr]. These approaches typically rely on table serialization, which is the process of converting table rows into text or sentences suitable for tokenization. For instance, [DBLP:conf/nips/0001PS24] fine-tuned a Llama 3-8B model on a large corpus of serialized tables and demonstrated that this strategy can outperform traditional tree-based models in few-shot scenarios. However, such language model–based approaches face inherent challenges. Their limited context windows restrict the number of serialized examples that can be processed simultaneously (e.g., up to 32 or 64 shots in [DBLP:conf/nips/0001PS24]), and it remains uncertain whether LLMs can reliably interpret and reason over numerical values [DBLP:conf/naacl/ThawaniPIS21].
Adopting a fundamentally different strategy, the authors of [tabpfn] introduced TabPFN, a transformer-based tabular foundation model designed for classification tasks and pretrained exclusively on synthetic tabular data. A key feature of TabPFN is its ability to perform in-context learning directly on tables, removing the need for tokenization and allowing efficient processing of relatively small datasets—up to 1K samples and 100 features. Building on this foundation, TabICL [tabicl] introduced a simplified three-component architecture comprising: (1) column-wise embeddings via Set Transformers to capture distribution-aware feature semantics, (2) row-wise interactions with rotary positional encodings to model inter-feature dependencies, and (3) dataset-level ICL prediction through split attention, ensuring a clear separation between training and test samples. These developments position tabular foundation models as a compelling alternative to traditional approaches, particularly for zero-shot prediction tasks where dataset-specific training is infeasible.
@@
The column-wise embedding component of Orion-MSP follows TabICL’s approach [tabicl], using Set Transformers with Induced Set Attention Blocks (ISAB) [isab] to create distribution-aware feature embeddings in a permutation-invariant manner. The multi-scale row interaction component processes these embeddings at multiple resolutions, with each scale using sparse attention patterns tailored to its granularity. The resulting multi-scale representations are aggregated into unified row embeddings, which then interact with the Perceiver memory before proceeding to the final ICL prediction stage. This ICL component employs split attention with label injection, ensuring proper train-test separation.
@@ 2  Related Work
Tabular In-Context Learning: The application of in-context learning (ICL) to tabular data has recently attracted significant attention. TabPFN [tabpfn] pioneered this direction by meta-training a transformer on synthetically generated datasets using structural causal models. Its encoder–decoder design allows test samples to attend to training examples, enabling zero-shot predictions without gradient-based fine-tuning. While TabPFN demonstrated strong performance on small datasets, its alternating column- and row-wise attention mechanisms make scaling to larger tables computationally prohibitive.
TabDPT [tabdpt] showed that comparable performance can be achieved on real-world datasets by using similarity-based retrieval to construct contextual examples—an idea first explored in TabR [tabr]. The authors extended this paradigm by integrating diffusion-based representation learning, improving robustness to missing values and distributional shifts. However, the diffusion process introduces substantial computational overhead and retains dense attention, limiting scalability. Similarly, TabPFN-v2 [tabpfn2] introduced cell-based in-context learning, extending row-wise encoding to datasets exceeding 10,00010{,}000 samples, but it still inherits quadratic attention costs in high-dimensional tables.
Building on these foundations, TabICL [tabicl] proposed a table-native transformer architecture with three components: column embedding via Set Transformers, row-wise interaction with rotary positional encodings, and an in-context learning prediction module. This design achieved state-of-the-art results across diverse benchmarks while maintaining architectural simplicity and training efficiency. Nonetheless, dense attention in row interactions and the strictly sequential pipeline limit iterative refinement, cross-component communication, and scalability to tables with more than 100 features.
ContextTab [contexttab] further enhanced tabular in-context learning by incorporating contextualized feature embeddings and attention mechanisms tailored for heterogeneous tabular data. While improving performance in complex datasets, it still processes features at a single scale and relies on dense attention, limiting computational efficiency on high-dimensional tables.
@@
Sparse Attention Mechanisms: Sparse attention techniques from natural language processing offer a promising route to improve computational efficiency in tabular in-context learning. BigBird [bigbird] and Longformer [longformer] demonstrated that block-sparse attention patterns can approximate dense attention with linear complexity while maintaining strong theoretical guarantees. Similarly, Sparse Transformers [sparse_transformers, sparse_transformers_neurips] employ structured sparsity for generative modeling, reducing computation without substantial performance degradation. Despite their success in sequential data, these methods have yet to be systematically adapted for tabular in-context learning, where the primary challenge lies in feature dimension rather than sequence length.
Hierarchical and Multi-Scale Architectures: Hierarchical architectures have proven effective in other domains. Funnel Transformers [funnel] and Swin Transformers [swin] use multi-scale processing and pooling to capture information at different resolutions, while Set Transformers [settransformers, settransformers2] leverage pooling by multihead attention for permutation-invariant set processing. Although TabICL [tabicl] employs Set Transformers for column embeddings, it does not incorporate hierarchical multi-scale processing or iterative pooling across feature groups, limiting its ability to model complex interactions in high-dimensional tables.
Cross-Component Communication: Cross-component memory and iterative refinement have shown success in multimodal learning. Perceiver [perceiver] and Perceiver IO [perceiverio] introduce latent bottlenecks to compress and share information across modalities, and vision-language models [vision_models] leverage iterative cross-attention for refinement. However, these approaches do not address the causal constraints of in-context learning, where test examples must never influence the representation of training data, leaving a gap for tabular in-context learning.
@@ 3.2  High-level Structure: From Data to ICL
Orion-MSP extends the original TabICL [tabicl] architecture with three complementary innovations designed to address the fundamental challenges of tabular data processing: computational inefficiency, limited feature interaction modeling, and the need for hierarchical pattern recognition. Our approach maintains the core in-context learning paradigm while introducing architectural enhancements that significantly improve both efficiency and performance.
@@ 3.3  Column-wise Embedding
Tabular data exhibits unique characteristics compared to other modalities: each column represents a distinct feature with its own distribution, scale, and statistical properties (e.g., mean, variance, skewness, kurtosis). To capture these distributional characteristics, we adopt the original TabICL [tabicl] column-wise embedder to map each scalar cell in a column $c_{j}\in\mathbb{R}^{n}$ to a $d$-dimensional representation using a shareable Set Transformer, $TF_{\text{col}}$, that treats the column as a permutation-invariant set of values. Our goal is to transform each cell value $X_{ij}$ into a $d$-dimensional embedding $\mathbf{E}_{ij}\in\mathbb{R}^{d}$ that encodes both:
@@ 3.3.2  Induced Set Attention Blocks (ISAB):
To efficiently capture global distributional information while maintaining computational tractability, we employ ISAB [isab] with $k$ learnable inducing points. It consists of two sequential Multi-Head Attention Blocks $(\text{MAB}_{1},\text{MAB}_{2})$:
@@
Following TabICL [tabicl], we use d=128d{=}128, k=128k{=}128, 4 heads, and 3 ISAB blocks. Crucially, in Equation 2, we use only training samples $\mathbf{U}_{j}^{\text{train}}\in\mathbb{R}^{n_{\text{train}}\times d}$ as keys and values. This ensures that the inducing points $\mathbf{M}_{j}$ capture the distribution of the training data only, preventing information leakage from test samples during embedding. This is crucial for maintaining the in-context learning paradigm.
@@ 3.4.1  Motivation and Design Principles
Inspired by hierarchical representations in vision [ying2018] and multi-resolution modeling in speech [2021wavesplit], Orion-MSP decomposes feature interactions into multiple resolution levels:
@@
To further improve efficiency and generalization, we adopt a block-sparse attention pattern inspired by Longformer [longformer] and BigBird [bigbird], as depicted in Figure 2:
@@ 3.4.2  Multi-Scale Feature Grouping:
The default grouping strategy uses a learnable soft grouping via Pooling by Multihead Attention (PMA) [pma] to adaptively attend to features:
@@ 3.4.5  Transformer Encoder per Scale:
  • Rotary positional encoding (RoPE): Applied to queries and keys before attention [roformer]
@@ 3.5  Cross-Component Memory with Perceiver Architecture
To overcome this limitation, we introduce a Perceiver-style latent memory module [perceiver] that enables safe, leak-free communication between architectural components. This latent memory acts as a shared representation space that can be written to by training samples and read from by both training and test samples, ensuring compliance with ICL constraints while promoting global knowledge sharing.
In standard transformer-based tabular architectures such as TabICL [tabicl], model components operate in a strictly sequential and isolated fashion:
@@
Inspired by the Perceiver architecture [perceiver], we introduce a learnable latent memory $\mathbf{L}\in\mathbb{R}^{P\times d}$ with $P$ memory slots. The key idea is:
@@ 4.1  Experimental Setting
Our experimental evaluation spans three widely recognized benchmark suites: TALENT [talent] (181 automatically discovered classification datasets), OpenML-CC18 [openmlcc18] (72 curated datasets), and TabZilla [tabzilla] (36 heterogeneous tasks). Together, these benchmarks enable a comprehensive assessment across diverse tabular learning scenarios. In addition, we perform domain-specific evaluations in high-impact application areas such as healthcare and finance to examine the real-world relevance of our method. All experiments strictly follow the official dataset splits provided by each benchmark to ensure reproducibility and fairness.
@@
Finally, we emphasize that models with higher mean ranks may not always achieve the highest absolute accuracy or F1-scores on every dataset. Rankings based on accuracy are computed per dataset and then averaged across all datasets, providing a normalized indicator of overall consistency rather than peak task-specific performance. In contrast, absolute metrics highlight maximum achievable performance on individual tasks. Comprehensive dataset statistics are presented in Appendix C.
@@
We compare our model with six state-of-the-art tabular foundation models: TabPFN [tabpfn], TabICL [tabicl], OrionBiX, Mitra, ContextTab [contexttab], and TabDPT [tabdpt]. In addition, we include established traditional baselines using autogloun [erickson2020autogluon] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.
@@
Performance. We measure predictive capability using standard classification metrics—Accuracy (ACC), AUC-ROC, and weighted F1-score (F1)—computed across the benchmark suites TALENT, OpenML-CC18, and TabZilla. These benchmarks encompass datasets with diverse characteristics, including varying sample sizes, feature dimensionalities, and class balance, allowing a comprehensive assessment of model generalization. It is important to clarify how mean rank values are derived. Within each benchmark suite, models are ranked by accuracy on every dataset (lower rank = better performance), and these per-dataset ranks are averaged to obtain the overall mean rank. Thus, a lower mean rank indicates stronger and more consistent performance across datasets, rather than the highest score on any single task. While absolute metrics (accuracy, F1) reflect peak task-level performance, mean rank provides a normalized measure of cross-dataset generalization consistency.
@@ 4.2  Results
Table 1: Performance comparison across three benchmark suites—TALENT, OpenML-CC18, and TabZilla. Ranks denote the mean rank based on accuracy per benchmark suite (lower is better). Metrics: ACC = Accuracy, F1 = Weighted F1. The “All” column reports the aggregated rank across all suites. Formatting: 1st place; 2nd place.
@@
Table 2: Performance variation by dataset size across all benchmark suites. Rank denotes the accuracy-based ranking per size category (lower is better). ACC = Accuracy; F1 = Weighted F1-score, averaged across datasets within each size category: Small (<<1K samples), Medium (1K–10K), and Large (>>10K).
@@
Rank ACC F1 Rank ACC F1 Rank ACC F1
XGBoost 7.70 0.8168 0.7964 6.88 0.8363 0.8314 5.41 0.8969 0.8920
CatBoost 7.88 0.8124 0.7935 6.47 0.8340 0.8264 5.48 0.8797 0.8733
Random Forest 8.55 0.7988 0.8187 7.16 0.8285 0.8221 7.30 0.8694 0.8628
LightGBM 7.80 0.8143 0.7789 6.94 0.8314 0.8226 5.63 0.8827 0.8764
TabICL 6.04 0.8301 0.8338 4.77 0.8486 0.8398 4.61 0.8802 0.8743
OrionBiX 6.32 0.8330 0.8150 5.48 0.8348 0.8260 4.42 0.8729 0.8670
OrionMSP 5.93 0.8232 0.8194 3.70 0.8494 0.8402 3.04 0.8843 0.8768
TabPFN 6.50 0.8325 0.8131 3.81 0.8557 0.8462 5.73 0.8783 0.8713
Mitra 13.88 0.4334 0.3236 11.59 0.3600 0.2553 11.11 0.3837 0.2754
ContextTab 9.60 0.7578 0.7363 9.52 0.6210 0.5566 10.22 0.6388 0.5638
TabDPT 5.48 0.8333 0.8271 5.40 0.8424 0.8339 5.26 0.8831 0.8765
@@
Table 3: Performance variation by feature dimensionality (dataset width) across all benchmark suites. Rank denotes the accuracy-based ranking averaged within each width category (lower is better). ACC = Accuracy; F1 = Weighted F1-score. Values are on a 0–1 scale (higher is better). Formatting: 1st place ; 2nd place within each group.
@@
Rank ACC F1 Rank ACC F1 Rank ACC F1
XGBoost 6.77 0.8222 0.8159 6.90 0.8482 0.8410 4.79 0.9140 0.9039
CatBoost 5.63 0.8145 0.8067 6.88 0.8441 0.8344 5.50 0.9157 0.9084
Random Forest 7.15 0.8005 0.7044 7.44 0.8410 0.8235 7.52 0.9034 0.8936
LightGBM 6.15 0.8128 0.7907 6.92 0.8458 0.8326 7.47 0.8999 0.8908
TabICL 5.14 0.8208 0.8119 4.61 0.8627 0.8549 6.46 0.9101 0.8936
OrionBiX 4.64 0.8112 0.8043 5.46 0.8510 0.8417 6.73 0.8859 0.8849
OrionMSP 3.76 0.8394 0.8314 4.09 0.8572 0.8478 5.69 0.8860 0.8837
TabPFN 5.30 0.8187 0.8092 4.07 0.8676 0.8589 6.141 0.9129 0.9111
Mitra 11.25 0.3737 0.2683 11.84 0.3886 0.2781 13.03 0.2521 0.1497
ContextTab 9.52 0.6391 0.5719 9.59 0.6480 0.5843 10.97 0.6017 0.5651
TabDPT 4.66 0.8262 0.8189 5.45 0.8566 0.8483 7.23 0.8845 0.8820